A firm with a WACC of 10% is considering the following mutually exclusive projects:

0 1 2 3 4 5
Project 1 -$500 $70 $70 $70 $215 $215
Project 2 -$400 $300 $300 $65 $65 $65

Which project would you recommend?

Select the correct answer.

a. Project 2, since the NPV2 > NPV1.
b. Neither Project 1 nor 2, since each project's NPV < 0.
c. Both Projects 1 and 2, since both projects have NPV's > 0.
d. Project 1, since the NPV1 > NPV2.
e. Both Projects 1 and 2, since both projects have IRR's > 0.
Answer #1

1:Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=70/1.1+70/1.1^2+70/1.1^3+215/1.1^4+215/1.1^5

=454.43

NPV=Present value of inflows-Present value of outflows

=454.43-500

=($45.57)(Approx)(Negative)

2:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=300/1.1+300/1.1^2+65/1.1^3+65/1.1^4+65/1.1^5

=654.25

NPV=Present value of inflows-Present value of outflows

=654.25-400

=$254.25(Approx)

Hence project 2 must be selected having higher and positive NPV.

Hence the correct option is:

a. Project 2, since the NPV2 > NPV1